OpenStudy (anonymous):

NaOH+HCL->NaCl+H2O If 10.0g NaOH are reacted with the stoichiometric amount HCL, how many moles of NaCl should be isolated? What would this weigh?

OpenStudy (anonymous):

Moles of NaOH=10/40=0.25 Moles of NaCl=moles of NaOH=0.25 Weight=0.25*58.5=14.625

OpenStudy (anonymous):

Thank you! I was halfway there and didn't even realize it.

OpenStudy (anonymous):

no problem:)